When the driver of a car travelling at 72km/h suddenly sees a pedestrian in front of the car , he applies the brakes. Consider the reaction time of the driver is 0.2s. if the maximum deceleration that the brakes can produce is 20m/s2, Find the maximum stopping distance for the car


1
Expert's answer
2022-06-01T12:48:52-0400
"d=d_1+d_2=vt+\frac{v^2}{2a}"

"d={\rm 20\: m/s*0.2\: s+(20\: m/s)^2/(2*20\: m/s^2)}\\
=14\:\rm m"